Q16) find the difference of each of the following signed 8-bit numbers and indicate which cause an arithmetic overflow to occur (a) 0010 1100 b -0101 1010 01011011 Де)! 111 1110 1011 1010 1101
Answer #1

Answer)

a) 0010 1100 - 0101 1010 = 0101110 (Decimal value equivalent to = - 46)

b) 0101 1011 - 1110 1011 = 10010000 (Decimal value equivalent to = -144)

c) 1110 1011 - 1010 1101 = 0111110 (Decimal value equivalent to = 62)

For the signed 8-bit numbers and their subtraction operations as above results, no overflow occur for any of them as they fit into the 8-bits perfectly.